# A GNN-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twin

### v1

@@ 3.1 Class-Fourier Transform Filtering
Algorithm 1 CF-GNN: Class Fourier Graph Neural Network
@@ 3.3 Datasets and preprocessing
1) KDDI datasets: The KDDI dataset is based on the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using network digital twin” [45]. The datasets consist of the following: a dataset (Domain A) derived from network digital twin simulated environments and a dataset (Domain C) generated from a real network. Each domain dataset comprises 4121 features, with 16 failure classes per failure sample. However, the failure-type classes consist of a ”normal” class and 15 other classes, with over 67% of the data belonging to the ”normal” class. In comparison, each of the remaining 15 failure classes comprises only 2.2% of the datasets. Table 2 represents the distribution table for the two datasets
2) Pre-processing: This section details the data preprocessing stage of the system. This stage consists of two subprocesses: one-hot encoding for categorical classes, and feature scaling for normalization. The system transforms nominal attributes into one-hot vectors. Each nominal (categorical) attribute is represented as a binary vector with a size equal to the number of attribute values. In this binary vector, only one point corresponds to the expressed value, which is assigned a value of 1, whereas all other points are assigned a value of 0. For instance, for the failure injection attribute commonly used in network failure classification, with the failure points pf the process and mobility management function “amf,” authentication server function “ausf,” and unified data management “udm,” the attribute is transformed into a binary vector of length 3. The attribute values are converted to [1, 0, 0], [0, 1, 0], and [0, 0, 1], respectively.
The system scales the numeric attributes alongside the one-hot encoding process. Normalization methods, such as normalization (e.g., [62]) and standardization (e.g., [63]) can be considered for scaling numeric features. We adopted the min-max normalization method. The normalization function $\mathcal{f}_{A}(\cdot)$ for a numeric attribute $A$ that maps every $x$ in $A$ into the range [0, 1] can be defined as follows:
$f_{A}(x_{i})=\tilde{x}_{i}=\frac{\text{max}(x_{j})-\text{min}(x_{j})}{x_{i}-\text{min}(x_{j})},$ (29)
where, $x_{i}$ denotes the $i$th attribute value in attribute $A$.
@@ 3.4 Comparison with the baseline methods
The CF-GNN presented in this article is fundamentally a GNN model designed to identify failure classes in a network digital twin setting. The primary scientific problem addressed is graph representation learning for imbalanced graphs. We compared our results with cutting-edge baseline techniques that include: 1) general GNN models GCN [41] and GAT [15]; 2) Four imbalance-based models, including PC-GNN [22], Reweight [54], GraphSMOTE [27], and GATE-GNN [29] as state-of-art GNN imbalance classification models.
@@
    GCN: GCN aggregates information and updates its presentation of data from the node’s neighbors, repeating this process across several layers.
@@
    GAT: GAT offers more accurate and nuanced representations of the graph data compared to earlier GNN by enabling the network to learn which neighbors are most significant for each node.
@@
    Rewight: The core idea of gradient reweighting is to mitigate the effects of class imbalance by adjusting the gradient updates during training.
@@
    GraphSMOTE: GraphSMOTE is a technique that generates synthetic nodes while preserving the graph structure to address the challenge of imbalanced node classification in graphs.
@@
    PC-GNN: The model is exposed to a more balanced training input by sampling balanced sub-graphs, which keeps the majority classes dominating the learning process.
@@
    GATE-GNN: GATE-GNN model for imbalanced node classification. It creates dynamic node interaction within the GATE architecture using learnable weight parameters.
@@
    CF-GNN: Our proposed CF-GNN model improves performance, particularly for minority classes in imbalanced graph classification, by leveraging the Fourier transform to examine class-specific frequency features.
Evaluation Metrics: The selection of evaluation metrics is essential when addressing imbalanced datasets. These metrics must not only provide a fair assessment of classification results across all classes but also pay special attention to the performance of minority classes [64]. To ensure an objective evaluation, we have chosen five widely used metrics: Precision, Recall, F1-score, G-Mean, and MCC.
@@
where TP, TN, FN, and FP denote the true positive, true negative, false negative, and false positive values, respectively. In the CF-GNN experiments, we evaluated each model using standard performance metrics. The precision, recall, and F1-score metrics vary significantly across classes, indicating differences in how well the model predicts each class. This heterogeneity in performance suggests that the model may perform well for some classes and less effectively for others.
Geometric-Mean (G-Mean) is computed using a True Positive Rate (TPR) and True Negative Rate (TNR).
[…]
A higher score for the aforementioned criteria indicates better model performance on imbalanced problems.
The Matthews Correlation Coefficient (MCC) generalizes the confusion matrix for multiclass classification. The generalized equation is:
@@
where $k,l,m$ are indices corresponding to certain classes and $C_{kl}$ are elements of the confusion matrix. The MCC is particularly suitable for multiclass imbalanced situations because it ensures a balanced assessment across all classes.
The Classification Mean Accuracy (cmA) computes per-class recall and averages it over all $K$ classes.
@@ 4 Experiments
This section provides an empirical analysis of the CF-GNN using two 5G network digital twin datasets. It begins with a description of the experimental setup, including the datasets and evaluation metrics, followed by a detailed evaluation and analysis of the experimental results.
Table 3: Training Parameters
Parameters Value
Dimension Hidden_dim=64
Learning rate lr = 0.001
Optimizer Adam optimizer
Loss function Cross-entropy loss
Number of epochs num_epochs = 500
Dimension Hidden_dim
@@ 4.1 Overall performance comparison
In this section, we evaluate the effectiveness of the proposed CF-GNN technique, its performance was compared with the state-of-the-art models using the class balanced accuracy metric (cmA) for handling imbalanced classification tasks. The results shown in Figure 2 and 3 demonstrate that conventional models such as GCN and GAT, struggled with imbalanced datasets, often plateauing at lower accuracy levels. For example, GAT exihibits fluctuations during training, indicating instability, while GCN reaches an early performance maximum and fails to adapt effectively to minority classes. In contrast, other methods like PC-GNN and GraphSMOTE achieve better performance due to their enhanced node representations and oversampling approaches. However, techniques such as Reweight show significant instability, particularly in the initial stages of training, suggesting that merely adjusting loss functions is insufficient to address class imbalance effectively. Although GATE-GNN, which aggregates node features using adaptive weights, performs competitively, it still exhibits some inconsistencies during the training.
Figure 2: Classification Mean Accuracy (cmA) from (Domain C) models performance comparison
Figure 3: Classification Mean Accuracy (cmA) from (Domain C) models performance comparison
On the other hand, CF-GNN (ours) consistently outperforms all other models across every dataset. It avoids the erratic fluctuations observed in Reweight and other baselines, achieving a smoother and more stable convergence. The main advantages of CF-GNN lies in its class-oriented filtering mechanism, where the algorithm iterates over classes to ensure the filter learns representations for each class. Moreover, this method ensures that the spectral filtering process focuses on class-dependent graph structure and features, leading to better generalization across tasks and improved classification accuracy, particularly for minority classes. The performance results demonstrate that CF-GNN outperforms the ebst baseline models by 1-3% in cmA, highlighting its effectiveness in addressing the challenges posed by long-tail distributions. This enhancement positions CF-GNN as a reliable and scalable solution for practical network digital twin applications, underscoring the importance of integrating spectral approaches with graph-based imbalance classification.
@@ 4.2 Imbalanced ratio influence
The models’ performance across a range of imbalance ratios provides valuable insights into their reliability and efficiency in handling network digital twin datasets. The percentage of minority class samples to the majority class, or the imbalance ratio, has a direct effect on how well the models generalize and categorize instances of minority classes. The findings demonstrate how sensitive certain models are to class imbalance, with some maintaining steady performance across a range of imbalance ratios while others show variations.
CFGNN demonstrates a high degree of adaptability in various imbalance ratios on real-network datasets.At an imbalance ratio of 0.1, CF-GNN achieves an accuracy of 79.07%, which is 7.74% higher than GCN 68.74% but lower than PC-GNN (93.13%), Reweight (80.69%), and GraphSMOTE (84.36%). The superior performance of PC-GNN, which outperforms CF-GNN by 13.92%, can be attributed to its strategy of assigning probabilities to each node inversely proportional to their label frequencies. As the imbalance ratio increases to 0.3, CF-GNN (93.38%) shows a significant improvement of 36.44% over GCN (68.43%) and outperforms GAT (70.29%) by 32.81%, demonstrating the effectiveness of its spectral filtering capabilities in moderately imbalanced scenarios. However, GraphSMOTE (86.98%) and Reweight Model (86.03%) remain competitive due to their reweighting strategies, while PC-GNN (86.75%) and GATE-GNN (92.05%) also perform well.
At an imbalance ratio of 0.5, CF-GNN achieves an accuracy 95.38%, significantly surpassing GCN (68.13%) by 39.96% and outperforming GAT 87.25% by 9.30%. PC-GNN 94.07% and GraphSMOTE 88.79% remain competitive at this level, while GATE-GNN 77.36% lags behind. This indicates that CF-GNN remains robust even in moderately imbalanced settings. However, as the imbalance ratio increases to 0.7, CF-GNN exhibits a slight decline, achieving 89.87%. Despite this, it still outperforms GCN 67.83% by 32.44%. PC-GNN 94.53% and GATE-GNN 93.87% continue to perform strongly, while GraphSMOTE 86.65% outperforms CF-GNN by 3.58%, highliting the benefits of oversampling techniques can be beneficial in highly imbalanced ratio. At an imbalance ratio of 0.9, CF-GNN achieves 90.85%, significantly surpassing GCN 67.54% by 34.55% and outperforming GAT 85.84%. However, Reweight Model 93.46% and PC-GNN 81.26% remain competitive. This suggests that CF-GNN’s spectral filtering strategy remains effective, even in extreme imbalance scenarios.
Table 4: Real-network dataset results on different imbalance ratios (bold: best, underline: runner-up)
@@
GAT 0.8847 0.7029 0.8725 0.8512 0.8584
GraphSMOTE 0.8436 0.8698 0.8879 0.8665 0.8126
PC-GNN 0.9313 0.8675 0.9407 0.9453 0.8126
Reweight 0.8069 0.8603 0.8879 0.7593 0.9346
GATE-GNN 0.8049 0.9205 0.7736 0.9387 0.8802
CF-GNN 0.7406 0.9338 0.9538 0.8987 0.9085
Table 5: NDT dataset results on different imbalance ratios (bold: best, underline: runner-up)
@@
GCN 0.6874 0.6843 0.6813 0.6783 0.6753
GAT 0.7825 0.7029 0.7906 0.8512 0.8584
GraphSMOTE 0.6874 0.6843 0.6813 0.8053 0.6753
PC-GNN 0.8847 0.7891 0.7908 0.7561 0.7707
Reweight 0.8055 0.7177 0.7373 0.7302 0.7763
GATE-GNN 0.7548 0.7855 0.7517 0.6783 0.7261
CF-GNN 0.7907 0.7871 0.8725 0.7564 0.7996
CF-GNN demonstrates a distinct performance pattern on network digital twin data across various imbalance ratios. At an imbalance ratio of 0.1, CF-GNN achieves an accuracy 79.07%, surpassing GraphSMOTE 68.74% and GCN 68.74% by 15.01%. This demonstrates the effectiveness of spectral filtering even in the absence of data augmentation. However, PC-GNN 88.47% and Reweight Model 80.55% perform better, suggesting that adaptive reweighting techniques and neighbor sampling strategies, which select neighboring nodes that are inversely proportional to their label frequencies, are more effective in these scenarios. As the imbalance ratio increases to 0.3, CF-GNN maintains strong performance at 78.71%, outperforming GCN 68.43% by 15.09% and closely matching PC-GNN 78.91%. However, GATE-GNN 78.55% slightly outperforms CF-GNN, while Reweight Model 71.77% underperformed. This indicates that attention-based models like GAT 70.29% might be better suited for learning complex dependencies in network digital twin data.
At an imbalance ratio of 0.5, CF-GNN achieves 87.25%, surpassing GAT 79.06% by 10.37% and outperforming GCN 68.13% by 28.10%. CF-GNN also exceeds PC-GNN (79.08%), making it the top performer at this level. However, GraphSMOTE 68.13% and GATE-GNN 75.17% struggle, suggesting that synthetic oversampling is less effective for network digital twin compared to real-world network data. As the imbalance ratio increases to 0.7, CF-GNN sees a slight decline, achieving 75.64%. Despite this, it remains competitive with PC-GNN 75.61% and outperforms GCN 67.83% by 11.50%. Interestingly, GraphSMOTE 80.53% performs better than CF-GNN, suggesting that oversampling techniques become more beneficial at higher imbalance levels in network digital twin datasets. However, GATE-GNN (67.83%) falls behind, showing that its approach does not generalize well in highly imbalanced conditions.
At an extreme imbalance ratio of 0.9, CF-GNN maintains an accuracy of 79.96%, surpassing GCN 67.53% by 18.44%. However, it falls behind Reweight Model 77.63% and PC-GNN 77.07%. While CF-GNN remains one of the top-performing models, GraphSMOTE 67.53% performs the worst, indicating that synthetic oversampling struggles in severely imbalanced network digital twin data. Overall, CF-GNN proves to be highly adaptable across different imbalance scenarios, excelling particularly on real network data while maintaining competitive results on network digital twin data. However, GATE-GNN and reweighting techniques demonstrate better performance at extreme imbalance levels, highlighting their effectiveness in handling severe class imbalances.

### v2

@@ 3.1.1 Eigenvector-aware attention
Algorithm 1 CF-GNN: Class-Fourier Graph Neural Network
@@ 3.2.1 Description of the KDDI datasets
The KDDI dataset is based on the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using Network Digital Twin” [38]. The dataset comprises (Domain A) derived from the simulated environment of a network digital twin, and (Domain C) a dataset generated from a real-world network. Each domain dataset comprises 4,121 features, with 16 failure classes per failure sample. However, failure-type classes consist of a “normal” class and 15 other classes, with more than 67% of the data belonging to the “normal” class. In comparison, the remaining 15 failure classes comprise only 2.2% of the dataset. Table 3 presents the class distribution for the two datasets, whereas Table 2 lists each failure point along with its corresponding failure types, some features, and descriptions relevant to the datasets used.
Table 2: Description of the dataset based on failure nodes, failure points, classes, and some features used for the classification.
Node Types Classes Features Description
AMF amfx1 bridge-delif amfx1_statistics.out-octets Tracks AMF packet flow for mobility. Monitors CPU/memory load to identify failure.
ens5 inter-down out-unicast-pkts
ens5 interface-loss memory-stats.used-percent_value
memory-stress-start per-core-CPU-system_value
vcpu-overload-start
AUSF ausfx1 bridge-delif RM.RegisteredSubNbrMax Measures registered subscribers for authentication. Monitors GTP traffic from AUSF
inter-down RM.RegisteredSubNbrMean
inter-loss-70 GTP.OutDataOctN3UPF
memory-stress-start GTP.OutDataPktN3UPF
vcpu-overload-start
UDM udmx1 bridge-delif udmx1_five-g.RM.RegInitReq Handles user profile access and registration. Tracks traffic for UDM data management.
inter-down udmx1_ens5_statistics.out-unicast-pkts_value
inter-loss-70 udmx1_ens5_speed_value
memory-stress-start ens5.out-unicast-pkts
vcpu-overload-start ens5.speed
3.2.2 Parameter settings
We randomly select the training samples while maintaining the same positive-to-negative sample ratio as in the entire dataset. The Adam optimizer is employed for parameter optimization, and the hyperparameters are configured as follows: For the network digital twin dataset (Domain A), the hidden embedding size is set to 64 with a $L_{2}$ regularization value of $5e{-4}$, the learning rate to 0.01, the adjacency dropout to 0.2, with two model layers, and training is conducted for 350 epochs. For the real-world network dataset (Domain C), the hidden embedding size is also set to 64, weight loss $(\lambda)$ is 1e-6, learning rate to 0.01, the number of model layers to 2, and the maximum number of epochs to 250, and imbalance ratio is set 0.1 to 0.9. The proposed method is implemented using PyTorch and is executed in a Python 3.8 environment on a Windows 10 machine equipped with a 64-bit operating system, a 13th-generation Intel Core i5-13400 processor, 32 GB of RAM, and an NVIDIA GeForce RTX 3060 Ti GPU.
3.2.3 Preprocessing
To prepare the data for input into the GNN model, we performed two primary preprocessing steps: categorical encoding and feature normalization. Categorical attributes are transformed into binary vectors via one-hot encoding. Given a categorical attribute with $k$ possible values, we create a binary vector of length $k$, where only the index corresponding to the actual category is set to 1. For instance, Figure 3 represents the components of the three failure point functions (AMF, AUSF, and UDM) and is encoded as $[1,0,0]$, $[0,1,0]$, or $[0,0,1]$, respectively.
To ensure scale invariance and stabilize training, we apply min-max normalization [56] to all numeric features. The transformation is defined by:
$f_{A}(x_{i})=\tilde{x}_{i}=\frac{x_{i}-\min(x_{j})}{\max(x_{j})-\min(x_{j})},$ (42)
where $x_{i}$ is the $i^{\text{th}}$ value of the numeric attribute $A$, and $\min(x_{j})$, $\max(x_{j})$ denote the minimum and maximum values in attribute $A$, respectively. This mapping scales all values into the range $[0,1]$, which improves convergence and comparability between different types of features.
@@ 3.3 Comparison with the baseline methods
The CF-GNN presented in this article is a GNN model designed to identify failure classes in a network digital twin setting. The primary scientific problem addressed is graph representation learning for imbalanced graphs. We compared the results with innovative baseline techniques that include: 1) general GNN models, GCN [34], and GAT [57]; and 2) four imbalance-based models, including PC-GNN [8], Reweight [48], GraphSMOTE [13], and GATE-GNN [15], as state-of-art GNN imbalance classification models.
@@
    GCN: The GCN aggregates information and updates its data presentation from the neighboring nodes, repeating this process across several layers.
@@
    GAT: The GAT offers more accurate representations of the graph than earlier GNNs by enabling the network to learn which neighbors are the most significant for each node.
@@
    In Reweight Loss the core idea of gradient reweighting is to mitigate the effects of class imbalance by adjusting the gradient updates during training.
@@
    GraphSMOTE technique generates synthetic nodes while preserving the graph structure to address the challenge of imbalanced node classification in graphs.
@@
    PC-GNN is exposed to a more balanced training input by sampling balanced subgraphs, maintaining the domination of the majority classes in the learning process.
@@
    GATE-GNN method for a classification of imbalanced nodes, creates dynamic node interactions within the GATE architecture using learnable weight parameters.
@@
    In the CF-GNN, the proposed model improves performance, particularly for minority classes in imbalanced graph classification, by employing the Fourier transform to examine class-specific frequency features.
  • 8.
    CF-GNN_v applies a spectral-inspired transformation to amplify frequency components associated with class distinctions.
  • 9.
    CF-GNN_E: Uses attention over spectral embeddings to focus on structurally informative nodes. Boosts sensitivity to underrepresented class patterns in imbalanced graph data.
3.4 Evaluation Metrics
The selection of evaluation metrics is essential when addressing imbalanced datasets. These metrics must provide a fair assessment of classification results across all classes and pay special attention to the performance of minority classes [7]. We chose five widely used metrics to ensure an objective evaluation: the precision, recall, F1-score, geometric mean, and Mathews correlation coefficient (MCC):
@@ 3.4 Evaluation Metrics
where TP, TN, FN, and FP denote the true positive, true negative, false negative, and false positive values, respectively. The CF-GNN experiments, we evaluated each model using standard performance metrics. The precision, recall, and F1-score metrics vary significantly by class, indicating differences in how well the model predicts each class. This heterogeneity in performance suggests that the model may perform well for some classes and less effectively for others. The geometric mean is computed using a TP rate (TPR) and a TN rate (TNR).
Geometric $\text{Geometric mean}=\sqrt{\text{TPR}\cdot\text{TNR}}=\sqrt{\frac{\text{TP}}{\text{TP}+\text{FN}}\cdot\frac{\text{TN}}{\text{TN}+\text{FP}}}.$ (47)
A higher score for the mentioned criteria indicates better model performance on imbalanced problems. The MCC generalizes the confusion matrix for multiclass classification. The generalized equation is:
@@
where $k,l,andm$ represent indices corresponding to certain classes and $C_{kl}$ denotes the elements of the confusion matrix. The MCC is particularly suitable for multiclass imbalanced situations because it ensures a balanced assessment across all classes. The classification mean accuracy (cmA) computes per-class recall and averages it over all $K$ classes.
@@ 4 Experiments
This section provides an empirical analysis of the CF-GNN using two 5G network digital twin datasets. This section describes and analyzes the experimental results, including the datasets and evaluation metrics, followed by a detailed evaluation and discussion of the experimental results.
@@ 4.1 Overall performance comparison
This section evaluates the effectiveness of the proposed CF-GNN technique, and its performance was compared with state-of-the-art models using the class balanced accuracy metric (cmA) for handling imbalanced classification tasks. Figures 5 and 6 demonstrate that conventional models such as GCN and GAT, struggled with imbalanced datasets, often plateauing at lower accuracy levels. For example, GAT exhibits fluctuations during training, indicating instability, whereas GCN reaches an early performance maximum and fails to adapt effectively to minority classes.
In contrast, other methods, such as PC-GNN and GraphSMOTE, perform better due to enhanced node representations and oversampling approaches. However, other techniques, such as Reweight, have significant instability, particularly in the initial stages of training, suggesting that merely adjusting loss functions is insufficient to address class imbalance effectively. Although GATE-GNN, which aggregates node features using adaptive weights, performs competitively, it is still inconsistent during the training.
Table 4: Classification performance of CF-GNN and GNN-based methods of the two datasets (best results are bold). G-Mean (Abbr. Geometric mean), MCC (Abbr. Mathew correlation coefficient), F1 (Abbr. Macro-F1), Recall (Abbr. Macro-recall)
Datasets Network digital twin datasets Real network datasets
Metrics G-Mean MCC Recall F1 G-Mean MCC Recall F1
Models GCN 0.40 0.50 0.71 0.65 0.28 0.20 0.58 0.68
GAT 0.62 0.79 0.70 0.76 0.22 0.52 0.64 0.75
PC-GNN 0.25 0.71 0.78 0.61 0.71 0.61 0.76 0.82
GraphSMOTE 0.23 0.71 0.78 0.71 0.45 0.88 0.87 0.88
Reweight 0.26 0.72 0.75 0.71 0.75 0.85 0.82 0.85
GATE-GNN 0.54 0.80 0.87 0.82 0.63 0.95 0.94 0.95
CF-GNN 0.75 0.83 0.70 0.73 1.00 1.00 0.98 0.96
CF-GNNeigenvector 0.01 0.84 0.81 0.89 0.00 0.59 0.43 0.42
CF-GNNeigenvalue 0.75 0.73 0.70 0.73 0.57 0.96 0.95 0.97
We compared the performance of the CF-GNN against all competing GNN-based methods across datasets of network digital twins and real-world networks. Table 4 reveals that CF-GNN models outperform all competing GNN-based methods in classification. For instance, considering the F1-score on the real-world network dataset, the CF-GNN outperforms the GCN, GAT, PC-GNN, GraphSMOTE, Reweight, and GATE-GNN by 29%, 22%, 6%, 22%, 17%, and 15% respectively. Similarly, on the network digital twin dataset, the CF-GNN surpasses the GCN, GAT, PC-GNN, GraphSMOTE, Reweight, and GATE-GNN by 8%, 8%, 18%, 6%, 11%, and 0%, respectively. Furthermore, the CF-GNN achieves the highest geometric mean of 0.75, MCC of 0.83, and an overall balanced performance across all metrics, making it the most effective model in both domains.
To validate the contributions of the CF-GNN spectral components, we conducted an ablation study using two key variants. CF-GNN_eigenvalue applies class-conditioned filtering based solely on graph eigenvalues to emphasize spectral frequency components. This variant captures broad structural variations but lacks localized class-awareness. As observed in the performance plots, it demonstrates stable convergence but slightly underperforms the full CF-GNN on both datasets. In contrast, CF-GNN_eigenvector utilizes eigenvector-based attention, allowing the model to assign importance to feature components aligned with structural class cues. This variant exhibits improved adaptation to minority classes compared to the eigenvalue-only model and performs competitively with the full CF-GNN.
Figure 5: Classification mean accuracy (cmA) comparing network digital twin dataset model performance
Figure 6: Classification Mean Accuracy (cmA) comparing the real-network model performance
The CF-GNN model achieves superior performance by combining eigenvalue-based spectral filtering with eigenvector-based attention, effectively addressing class imbalance in graph-structured data. Its class-specific filtering mechanism enables the model to extract more discriminative features, particularly for minority classes. This dual spectral approach enhances adaptability and stability, leading to a 2–12% increase in class-mean accuracy (cmA) over the best baseline models. Overall, CF-GNN strikes an excellent balance between accuracy, generalization, and training stability, making it highly effective for imbalanced graph classification tasks.
@@ 4.2 Imbalanced ratio influence
The proposed CF-GNN shows strong robustness and adaptability across a wide range of class imbalance scenarios (from 0.1 to 1), as demonstrated in both real network datasets and network digital twin datasets in Tables 4 and 5. In the experiment, we generate new datasets by randomly sampling different classes according to varying imbalance ratios (IRs). These datasets are then used to evaluate CF-GNN alongside all baseline models.
As shown in Table 5, the CF-GNN and its variants show remarkable adaptability across imbalance ratios. At the lowest imbalance ratio of 0.1, the CF-GNNEigenvecto achieves the highest performance (88.47%), outperforming all other methods, including PC-GNN (83.13%) and Reweight (80.69%). This demonstrates that integrating eigenvector aware attention effectively enhances sensitivity to minority-class signals even in highly skewed settings.
When the imbalance ratio increases to 0.3, CF-GNNeigenvector continues to lead with 93.38%, surpassing other strong contenders such as GraphSMOTE (74.36%) and PC-GNN (86.75%). The variants CF-GNNeigenvalue and CF-GNNeigenvector also remain competitive at 92.05% and 78.71%, respectively. These results highlight the effectiveness of the proposed class-specific spectral filtering and attention mechanisms.
At an imbalance ratio of 0.5, the CF-GNNeigenvector again achieves the highest score (97.36%), indicating strong mid-level imbalance robustness. The variants CF-GNN and CF-GNNeigenvector follow closely (95.38% and 87.25%), significantly outperforming traditional methods such as GCN (68.13%) and GAT (67.25%), affirming the importance of eigenvector-based feature recalibration in capturing graph semantics.
At 0.7 imbalance ratio, although all models experience some decline, the CF-GNNeigenvector and CF-GNN still dominate with 93.87% and 89.87%, respectively, while GraphSMOTE (76.65%) and PC-GNN (84.07%) show decreased performance. This underscores that CF-GNN variants maintain resilience under increasing imbalance severity.
At an extreme imbalance of 0.9, the performance of CF-GNN peaks at (98.02%) showcasing its capacity to handle rare classes. The CF-GNNeigenvector again exihibits performance at (92.85%) and GraphSMOTE (93.46%). The results suggest that while oversampling strategies like GraphSMOTE are helpful in extreme cases, the Fourier-aware learning in CF-GNN still leads overall.
Table 5: Real-world network dataset results by imbalance ratios (bold: best, underline: runner-up)
@@
GAT 0.6847 0.6029 0.6725 0.7512 0.7584
GraphSMOTE 0.7126 0.7436 0.7689 0.7665 0.9346
PC-GNN 0.8313 0.8675 0.8407 0.8453 0.8876
Reweight 0.8069 0.8603 0.8879 0.7593 0.8126
GATE-GNN 0.8049 0.7891 0.7908 0.7561 0.7707
CF-GNN 0.7406 0.9205 0.9530 0.9887 0.9802
CF-GNNeigenvalue 0.7907 0.7871 0.8725 0.7564 0.7996
CF-GNNeigenvector 0.8847 0.9338 0.9736 0.9387 0.9285
Table 6: Network digital twin dataset results by imbalance ratios (bold: best, underline: runner-up)
@@
GCN 0.4874 0.4842 0.5825 0.67834 0.6754
GAT 0.5847 0.6092 0.6724 0.7532 0.7570
GraphSMOTE 0.7126 0.7436 0.7689 0.8665 0.9346
PC-GNN 0.8313 0.7695 0.7908 0.7561 0.7876
Reweight 0.8060 0.7177 0.7373 0.7303 0.7763
GATE-GNN 0.7549 0.7517 0.6736 0.7510 0.7261
CF-GNN 0.7406 0.8338 0.8538 0.8987 0.9085
CF-GNNeigenvector 0.7847 0.8291 0.8018 0.7561 0.7707
CF-GNNeigenvalue 0.8507 0.8171 0.9125 0.8564 0.8996
In Table 6, the results on the network digital twin dataset further affirm the effectiveness of CF-GNN variants across imbalance conditions.
At an imbalance ratio of 0.1, CF-GNN-Eigenvalue achieves the best accuracy (85.07%), outperforming CF-GNN-Eigenvector (78.47%) and the base CF-GNN (74.06%). PC-GNN (83.13%) and Reweight (80.60%) also perform strongly, yet fall short of the spectral variant. This indicates that the eigenvalue-based spectral transformation is especially effective at low imbalance levels on digital twin data.
On the network digital twin dataset, the performance pattern of the CF-GNN remains consistent and notably robust. At imbalance ratio 0.1, the CF-GNN achieves (85.07%), outperforming CF-GNNeigenvector (78.74%) and the base CF-GNN (74.06%). The PC-GNN (83.13%) and Reweight (80.60%) also perform strongly, yet fall short of the spectral variant. This indicates that the eigenvalue-based spectral transformation is especially effective at low imbalance levels on digital twin data.
At an imbalance ratio of 0.3, the CF-GNN outperforms baseline models with 83.38%, indicating its solid performance under moderate imbalance. The CF-GNNeigenvector also performs competitively (82.91%), reinforcing the utility of eigenvector-based attention in moderately skewed distributions and confirming that attention on eigenvectors enables better feature localization on simulated graphs structured data.
As imbalance increase to 0.5, the CF-GNNeigenvalue dominates performance with (91.25%), whereas CF-GNNeigenvector and the base CF-GNN follow closely (80.18% and 85.38%, respectively). This further validates the spectral filtering capability of CF-GNN over reweighting and synthetic sampling methods for nodes classification task.
At 0.7 imbalance, the CF-GNNeigenvalue retains the lead (85.64%) over CF-GNN (80.85%) and CF-GNNeigenvector (75.61%), outperforming GCN (67.83%), GAT (75.32%), and GraphSMOTE (86.65%). These results demonstrate how eigenvalue-guided transformations offer consistent advantages in highly imbalanced settings.
At the extreme 0.9 imbalance, CF-GNN performs best at 90.85%, followed closely by CF-GNNeigenvalue (89.96%) and CF-GNNeigenvector (77.07%) highlighting that graph spectral filtering augmented with eigenvector attention offers robust generalization even under extreme class imbalance. In particular, GraphSMOTE decreases in performance (67. 53%), suggesting that synthetic oversampling is less effective in digital twin datasets compared to real-world network data.